Flamo:

Find the length of the hypotenuse of an isosceles right triangle whose legs are 1 unit in length.

Flamo:

@Hero, I need help with this Fast.

Flamo:

I don't know what to do.

Angle:



Angle:

this is the triangle we're looking at

Flamo:

o, Ye

Angle:

have you learned the pythagorean theorem yet?

Flamo:

Nope.

Flamo:

Isn't it a^2 + b^2 = c^2

Flamo:

The Formula?

Angle:

exactly

Flamo:

o

Angle:

ok so a and b are the legs and c is the hypotenus

Flamo:

o

Angle:

so a = 1 b = 1 c = ? plug into the equation

Flamo:

1^2 + 1^2 = c^2

Flamo:

?

Angle:

perfect!

Flamo:

1^2 = 1, So 1 + 1 = c^2

Angle:

mhm

Flamo:

2 = c^2?

Angle:

yes but we want "c"

Angle:

so we should square root both sides because square root of (c^2) = c

Flamo:

o

Flamo:

So the Square root of 2 = 1.41

Flamo:

So the answer is 1.41?

Angle:

yes, but I think you can just leave your answer as "square root of 2"

Flamo:

o, Okay, Ty!